Today the team was able to start wrapping up deployment.
We were able to get the following completed:
- NFS fixed, tuned and mounted on cpus
- Lustre and client rpms were prepared and be ready tomorrow morning
- Configured provisional external ip addresses on the head nodes as well as login nodes.
- Tests were run to for GPFS NFS mount
- Updates to firmware were done and reapplied BIOS settings to the machines that had motherboards replaced due to the power outage yesterday.
- Started Linpack tests on another rack pair of rack. This is the last pair!
A Dell technician replaces the mother board on a node that was damaged during the power outage.